OpenStudy (anonymous):

Suppose x coins are tossed. Write an expression to represent the number of possible outcomes.

OpenStudy (anonymous):

two possible outcomes H or T for each toss, counting principle tells you if you do this \(x\) times there are \[2\times 2\times 2\times ...\times 2=2^x\] outcomes

OpenStudy (anonymous):

If 1 coin is tossed there would be 2 possible outcomes, (Head or Tail). Notice that the number of ways is 2 or 21. If 2 coins are tossed there are 2 possible outcomes for the 1st coin. Then for each of those 2 possible outcomes for the 1st coin, there are 2 possible otcomes for the 2nd coin. That's 2·2 or 4 ways. Notice that the number of ways for 2 coins is 2·2 or 22. If 3 coins are tossed there are 2 possible outcomes for the 1st coin. Then for each of those 2 possible outcomes for the 1st coin, there are 2 possible otcomes for the 2nd coin. That's 2·2 or 4 ways for the first two coins. Then for each of those 4 possible outcomes for the 1st coin, there are 2 possible otcomes for the 3rd coin. That's 2·2·2 or 4·2 or 8 ways for the three coins. Notice that the number of ways for 3 coins is 2·2·2 or 23. Since 1. The number of ways for 1 coin is 21 and 2. The number of ways for 2 coins is 22 and 3. The number of ways for 3 coins is 23. 4. ... So we can generalize and say: x. The number of ways for x coins is 2x.

OpenStudy (anonymous):

what @elyise said 'cept he\she means \(2^x\) not \(2x\)

OpenStudy (anonymous):

im a he and yes i mean \[2^{x}\]